40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 100     Entry time: Thu Oct 9 08:47:54 2008
Author: Aidan 
Type: Laser 
Category: Fiber 
Subject: Dismantled fiber stabilization setup 


Packed up the current setup. Am going to rebuild with beam double passed through the fiber and an 80MHz AOM.
ELOG V3.1.3-